# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

### v1

@@ Abstract
Among guessing random additive noise decoding (GRAND) algorithms, ordered reliability bits GRAND (ORBGRAND) has attracted considerable attention due to its efficient use of soft information and suitability for hardware implementation. It has also been shown that ORBGRAND achieves a rate very close to the capacity over additive white Gaussian noise channels with antipodal inputs. In this work, it is further established that, via suitably companding the ranks in ORBGRAND according to the inverse cumulative distribution function (CDF) of channel reliability, the resulting CDF-ORBGRAND algorithm exactly achieves the mutual information of general binary-input memoryless channels under symmetric input distribution, i.e., the symmetric capacity. This result is then applied to bit-interleaved coded modulation (BICM) systems to handle high-order input constellations. Via considering the effects of mismatched decoding due to both BICM and ORBGRAND, it is shown that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived by treating BICM as a set of independent parallel channels.
@@ I Introduction
Guessing random additive noise decoding (GRAND) [1, 2] has recently attracted attention as a new approach to decoding in next-generation communication scenarios. GRAND operates by progressively guessing the error pattern (EP) that the transmitted codeword has experienced through the channel. This approach centering around EPs makes the GRAND paradigm well suited for codes in the high-rate high-reliability regime.
@@
GRAND was originally designed for hard-decision channels. This is also reflected in its name: only when the noise is discrete it is meaningful to “guess” it. For more general channels, such as additive white Gaussian noise (AWGN) channels, incorporating channel soft information can improve decoding performance [3, Ch. 10], and we need to guess the EP rather than the noise itself. Symbol reliability GRAND (SRGRAND) [4] uses a single-bit reliability indicator to mark whether a channel output symbol is reliable. Soft GRAND (SGRAND)[5] directly exploits the exact log-likelihood ratio (LLR) magnitudes of the channel output to determine the sequence of EPs for guessing, thereby fully leveraging the soft information and achieving maximum-likelihood (ML) decoding. However, the requirement of exact LLR magnitudes in SGRAND necessitates on-the-fly generation of EPs, posing challenges for hardware implementation. Another variant of GRAND, ordered reliability bits GRAND (ORBGRAND) [6] generates EPs based on the ranking of LLR magnitudes rather than their exact values. This approach enables offline generation of EPs and streamlines hardware implementation [7, 8, 9]. It has also been shown in [10] through an information-theoretic analysis that for AWGN channels with antipodal inputs, the rate achieved by ORBGRAND is very close to the channel mutual information, i.e., the capacity in that setup. There also exist several works that treat GRAND for more general channels such as fading channels [11, 12, 13, 14, 15, 16].
@@
In this paper, we consider a modification of ORBGRAND, which essentially adopts a companding technique to transform the rank of each LLR magnitude according to its inverse cumulative distribution function (CDF). The resulting algorithm is termed CDF-ORBGRAND, whose basic idea has initially appeared in [6] and formal description provided in [17, 18]. Via such a companding technique, the decoding metric is rendered more aligned with the channel law, while still maintaining its amenability to hardware implementation. We analyze the achievable rate of CDF-ORBGRAND for general binary-input memoryless channels under symmetric input distribution, and prove via a random coding argument akin to that in [10] that, CDF-ORBGRAND exactly achieves the mutual information, i.e., the symmetric capacity, without any rate loss. In contrast, for fading channels, ORBGRAND without rank companding may incur a significant rate loss compared to the symmetric capacity; see Fig. 1 in Section III-A.
@@
On the basis of CDF-ORBGRAND for binary-input channels, we proceed to the investigation of bit-interleaved coded modulation (BICM) systems [19, 20]. This investigation is motivated by the prevalence of high-order modulation in modern communication systems, wherein BICM has been an extensively used technique in fading channels. By jointly considering BICM and GRAND as a mismatched decoding rule, we derive achievable rates of both CDF-ORBGRAND and ORBGRAND for BICM. Remarkably, we show that CDF-ORBGRAND achieves the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels [19], and subsequently established by typicality decoding or by mismatched decoding with appropriately chosen metrics [21]. The BICM capacity-achieving property of CDF-ORBGRAND indicates that the ranking companding technique is effective even for handling the heterogeneous bit channels corresponding to different positions of the labeling of constellation symbols. In contrast, without rank companding, ORBGRAND incurs a substantial rate loss which becomes even more pronounced as modulation order increases, as confirmed by numerical results.
@@
The remaining part of this paper is organized as follows: Section II introduces the system model, along with GRAND and its variants for general binary-input memoryless channels. Section III proves that CDF-ORBGRAND exactly achieves the symmetric capacity. Section IV considers a BICM system model, and applies GRAND therein. It further proves that CDF-ORBGRAND achieves the BICM capacity. Section V concludes this paper. Technical proofs are provided in Appendices.
@@ II-A System Model
Our basic channel model is a general binary-input memoryless channel, which, without loss of generality, is of input alphabet $\{+1,-1\}$, with the output probability density function (pdf) denoted by $q^{+}(y)$ when input $x=+1$, and by $q^{-}(y)$ when input x=1x=-1. Here $q^{+}(y)$ and $q^{-}(y)$ are general, without assuming any symmetry.
@@
and call their magnitudes, $\lvert\mathsf{T}_{1}\rvert,\cdots,\lvert\mathsf{T}_{N}\rvert$], the channel reliability vector. The CDF of $\lvert\mathsf{T}\rvert$ is denoted by $\Psi(t)$, for $t\geq 0$. We also introduce the sets $\mathcal{R}_{1}=\left\{y\lvert q^{+}(y)<q^{-}(y)\right\}$ and $\mathcal{R}_{2}=\left\{y\lvert q^{+}(y)>q^{-}(y)\right\}$.
@@
For technical reasons, we impose a few additional assumptions on the behavior of $q^{+}(y)$ and $q^{-}(y)$, as follows.
@@ II-B GRAND and Its Variants
In general, we can collectively denote the EPs used by GRAND as a $2^{N}\times N$ $\pm 1$-valued matrix $P$: in the $q$-th query, the $q$-th row of $P$ is used as the EP for testing, such that if $P_{q,i}=-1$ the sign of the $i$-th entry of $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output; otherwise, if all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared.
For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i:P_{q,i}=-1}r_{i}$, is non-decreasing with $q$, where $r_{i}$ is the rank of $\lvert t_{i}\rvert$ introduced in the previous subsection. Efficient algorithms for generating such $P$ are available (see, e.g., [6, 7]). Since $2^{N}$ is typically an exceedingly large quantity, in practice $P$ is usually truncated to its first $Q$ rows, with $Q$ being the maximum number of queries allowed.
As shown in [10, Sec. II], ORBGRAND and many other variants of GRAND can be represented by the following unified decoding rule, as long as no truncation of $P$ is applied, i.e., $Q=2^{N}$:
@@
The form of (2) facilitates information-theoretic analysis, as will be conducted in the following sections. A decoding error event occurs if $\hat{\mathsf{W}}\neq\mathsf{W}$, and we define the achievable rate of GRAND as follows.
@@
If we do not restrict the decoding rule to be (2) and apply ML decoding, then according to Shannon’s channel coding theorem, the supremum of all achievable rates is the mutual information $I(\mathsf{X};\mathsf{Y})$ with $\mathsf{X}$ uniformly distributed over $\{+1,-1\}$. This mutual information thus serves as the ultimate performance limit for the considered system model, and is termed the symmetric capacity in the sequel.
@@ III Capacity-achieving Property of CDF-ORBGRAND
To motivate CDF-ORBGRAND, it is helpful to inspect the decoding rule (2). Based on the convergence of empirical distributions, as $N$ gets large, the distribution of $\mathsf{R}_{i}/(N+1)$ gets close to that of $|\mathsf{T}_{i}|$, i.e., $\Psi$. [^2] This observation has been used in [10] to explain the behavior of ORBGRAND, by introducing a variant of GRAND termed cdf-GRAND therein. Unfortunately, cdf-GRAND relies on the exact channel reliability vector, and hence is not amenable to hardware implementation. Thus, by conducting the inverse operation, i.e., $\Psi^{-1}$, over $\mathsf{R}_{i}/(N+1)$, we essentially return to $|\mathsf{T}_{i}|$, and this is exactly what CDF-ORBGRAND does. Due to this intuitive argument, CDF-ORBGRAND should behave similarly to SGRAND, and this assertion will be formalized and established in this section. Note that in CDF-ORBGRAND since $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$ simply compands the rankings $\left\{\mathsf{R}_{i}\right\}_{i=1,\cdots,N}$, which can be computed and stored offline, its implementation is still similar to that of ORBGRAND, amenable to parallelism [22].
@@ III-A Main Result
Neither ORBGRAND nor CDF-ORBGRAND is ML decoding, and furthermore, their metrics, $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$, are $N$ correlated random variables due to ranking, we cannot directly invoke existing information-theoretic formulas for evaluating their achievable rates. In this paper, we essentially establish the achievable rate of CDF-ORBGRAND as the generalized mutual information (GMI) of the considered decoding rule, following the approach of mismatched decoding (see, e.g., [23, 24]). As revealed by our result below, although CDF-ORBGRAND is inherently mismatched, its achievable rate surprisingly coincides with the symmetric capacity. In other words, from an information-theoretic perspective, no rate loss is incurred when applying CDF-ORBGRAND.
@@
In our previous work[25], we derived the GMI of ORBGRAND for the system model in Section II-A as
@@
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Psi\Bigl(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\Bigr)q^{+}(y)\mathrm{d}y$
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Psi\Bigl(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\Bigr)q^{-}(y)\mathrm{d}y\Bigg\}$ (4)
@@
We conduct some numerical experiments to compare (4) and (5), for antipodal modulation over AWGN, additive white Laplacian noise (AWLN), and Rayleigh fading channels with receiver-side perfect channel state information (CSI). These channels can be written as
$\mathsf{Y}=\sqrt{P}\mathsf{H}\mathsf{X}+\mathsf{Z},$
where $\mathsf{H}$ is the channel coefficient and $\mathsf{Z}$ is the additive noise. The corresponding LLR $\mathsf{T}$ is:
@@
  • Rayleigh fading channel with receiver-side perfect CSI, $\mathsf{H}\sim\mathcal{CN}(0,1)$ and $\mathsf{Z}\sim\mathcal{CN}(0,1)$: $\mathsf{T}=4\sqrt{P}\Re(\mathsf{H}^{*}\mathsf{Y})$. We point out that here $q^{\pm}(\cdot)$ should be evaluated for the augmented channel output $(y,h)$, i.e., including the receiver-side perfect CSI. [^4] We clarify that in the numerical experiments in [25], the GMI of ORBGRAND for Rayleigh fading channels with receiver-side perfect CSI was incorrectly computed, due to an error in the program evaluating the integrals in (4). In the current paper, we rectify the error and provide updated numerical results.
@@ III-B Proof of Main Result
Since the terms in the summation in (6) involve rankings of channel reliability vector, they are correlated, and hence the standard GMI formula (see, e.g., [23, Eqn. (12)]) cannot be directly applied. Therefore we conduct the analysis directly from the first principle, akin to that in [10, 25]. We calculate the ensemble average error probability of decoding, which, for the i.i.d. random coding ensemble in our system model, is equivalent to the error probability of decoding when message $w=1$ is transmitted.
Under the condition that the transmitted message is $w=1$, some asymptotic properties of the CDF-ORBGRAND decoding metric (6) are derived and presented in the following lemmas. Their proofs are provided in the Appendix.
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=$ $\mathbb{E}\Bigl[\ln\Bigl(1+e^{\theta\cdot\left\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\right\rvert}\Bigr)\Bigr]-\ln 2.$ (9)
@@
$=\quad$ $\text{Pr}[\hat{\mathsf{W}}\neq 1\lvert\mathcal{U}_{\epsilon}]\text{Pr}[\mathcal{U}_{\epsilon}]+\text{Pr}[\hat{\mathsf{W}}\neq 1,\mathcal{U}_{\epsilon}^{c}]$
@@
$\lim\limits_{N\to\infty}$ $\text{Pr}\Bigg[\mathsf{D}(1)\geq\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]=0.$ (11)
@@
On the other hand, applying the decoding rule (2) together with the union bound, we obtain
@@
$\leq\quad$ $\text{Pr}\Bigg[\exists w^{\prime}\neq 1,\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]$
@@
Considering the probability in (12) conditioned with respect to $\underline{\mathsf{T}}$, and applying Chernoff bound, we obtain that for any $N$ and any $\theta<0$,
$-\frac{1}{N}\ln\text{Pr}\Bigg[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Big\lvert\underline{\mathsf{T}}\Bigg]$
$\geq\quad$ $\theta\Bigg[\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]$
$-\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}.$ (13)
@@
$\text{Pr}\Bigg[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Big\lvert\underline{\mathsf{T}}\Bigg]$
@@
$\hskip 40.00006pt+\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigg]\Bigg\}.$ (14)

### v2

@@ Abstract
Among guessing random additive noise decoding (GRAND) algorithms, ordered reliability bits GRAND (ORBGRAND) has attracted considerable attention due to its efficient use of soft information and suitability for hardware implementation. It has also been shown that ORBGRAND achieves a rate very close to the capacity of an additive white Gaussian noise channel under antipodal signaling. In this work, it is further established that, for general binary-input memoryless channels under symmetric input distribution, via suitably companding the ranks in ORBGRAND according to the inverse cumulative distribution function (CDF) of channel reliability, the resulting CDF-ORBGRAND algorithm exactly achieves the mutual information, i.e., the symmetric capacity. This result is then applied to bit-interleaved coded modulation (BICM) systems to handle high-order input constellations. Via considering the effects of mismatched decoding due to both BICM and ORBGRAND, it is shown that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels.
@@ I Introduction
Guessing random additive noise decoding (GRAND) [1, 2] has recently attracted attention as a new paradigm of decoding in emerging communication scenarios. GRAND operates by progressively guessing the error pattern (EP) that the transmitted codeword has experienced through the channel. This approach centering around EPs makes the GRAND paradigm well suited for channel coding in the high-rate high-reliability regime [3].
@@
GRAND was originally proposed for hard-decision channels. This is also reflected by its name: only when the noise is discrete it makes sense to “guess” it. For more general channels, such as additive white Gaussian noise (AWGN) channels, incorporating channel soft information can improve decoding performance [4, Ch. 10], and we need to guess the EP rather than the noise itself. Symbol reliability GRAND (SRGRAND) [5] uses a single-bit reliability indicator to mark whether a channel output symbol is reliable. Soft GRAND (SGRAND)[6] directly exploits the exact log-likelihood ratio (LLR) magnitudes of the channel output to determine the sequence of EPs for guessing, thereby fully leveraging the soft information and achieving maximum-likelihood (ML) decoding. However, the requirement of exact LLR magnitudes in SGRAND necessitates on-the-fly generation of EPs, posing challenges for hardware implementation. Another variant of GRAND, ordered reliability bits GRAND (ORBGRAND) [7] generates EPs based on the ranking of LLR magnitudes rather than their exact values. This approach enables efficient generation of EPs without depending upon exact LLR magnitudes, and streamlines hardware implementation [8, 9, 10]. It has also been shown in [11] through an information-theoretic analysis that for AWGN channels under antipodal signaling, the rate achieved by ORBGRAND is very close to the channel mutual information, i.e., the capacity in that setup. There also exist several works that treat GRAND for more general channels such as fading channels [12, 13, 14, 15, 16, 17].
@@
In this paper, we study a modification of ORBGRAND, which essentially adopts a companding technique to transform the rank of each LLR magnitude according to the inverse cumulative distribution function (CDF) of LLR magnitude. The resulting algorithm is thus termed CDF-ORBGRAND. The value of the shape of the CDF of LLR magnitude has been recognized earlier in [7], and a piecewise linear approximation technique has been utilized therein under the term of “full ORBGRAND”; see also [18, 19]. Via such a companding technique, the decoding metric is rendered asymptotically aligned with the channel law, while still maintaining its amenability to hardware implementation. We analyze the achievable rate of CDF-ORBGRAND for general binary-input memoryless channels under symmetric input distribution, and prove via a random coding argument akin to that in [11] that, CDF-ORBGRAND exactly achieves the mutual information, i.e., the symmetric capacity, without any rate loss. In contrast, for fading channels, ORBGRAND without rank companding may incur a significant rate loss compared to the symmetric capacity; see Fig. 1 in Section III-A.
@@
On the basis of CDF-ORBGRAND for binary-input channels, we proceed to the investigation of bit-interleaved coded modulation (BICM) systems [20, 21]. This investigation is motivated by the prevalence of high-order modulation in modern communication systems, wherein BICM has been a widely used technique. By jointly considering the mismatched decoding effects due to both BICM and ORBGRAND, we derive achievable rates of both CDF-ORBGRAND and ORBGRAND for BICM. Remarkably, we show that CDF-ORBGRAND achieves the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels [20], and subsequently established by typicality decoding or by mismatched decoding with appropriately chosen metrics [22]. The BICM capacity-achieving property of CDF-ORBGRAND indicates that the rank companding technique is effective even for handling the heterogeneity among bit channels corresponding to different positions of the labeling of constellation symbols. In contrast, without rank companding, the rate loss of ORBGRAND is evident and becomes more substantial as modulation order increases; see Fig. 4 in Section IV-E.
@@
The remaining part of this paper is organized as follows: Section II introduces the system model, along with GRAND and its variants for general binary-input memoryless channels. Section III proves that CDF-ORBGRAND exactly achieves the symmetric capacity. Section IV considers applying ORBGRAND and CDF-ORBGRAND to a BICM system model, and derives their respective achievable rates, revealing that CDF-ORBGRAND achieves the BICM capacity. Section V concludes this paper. Technical proofs are provided in Appendices.
@@ II-A System Model
Our basic channel model is a general binary-input memoryless channel, which, without loss of generality, is of input alphabet $\{+1,-1\}$, with the output probability density function (pdf) denoted by $q^{+}(y)$ when input $x=+1$, and by $q^{-}(y)$ when input x=1x=-1. Here $q^{+}(y)$ and $q^{-}(y)$ are general, without assuming any specific structure like symmetry.
@@
and call their magnitudes, $\lvert\mathsf{T}_{1}\rvert,\cdots,\lvert\mathsf{T}_{N}\rvert$], the channel reliability vector. The CDF of $\lvert\mathsf{T}\rvert$ is denoted by $\Psi(t)$, for $t\geq 0$. We also introduce the sets $\mathcal{R}_{1}=\left\{y|q^{+}(y)<q^{-}(y)\right\}$ and $\mathcal{R}_{2}=\left\{y|q^{+}(y)>q^{-}(y)\right\}$.
@@
For our analysis, we impose a few additional assumptions on the behavior of $q^{+}(y)$ and $q^{-}(y)$, as follows.
@@ II-B GRAND and Its Variants
In general, we can collectively denote the EPs used by GRAND as a $2^{N}\times N$ $\pm 1$-valued matrix $P$: in the $q$-th query, the $q$-th row of $P$ is used as the EP for testing, such that if $P_{q,i}=-1$ the sign of the $i$-th entry of $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output. If all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared.
For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i:P_{q,i}=-1}r_{i}$, is non-decreasing with $q$, where $r_{i}$ is the rank of $\lvert t_{i}\rvert$ introduced in the previous subsection. Efficient algorithms for generating such $P$ are available (see, e.g., [7, 8]). Since $2^{N}$ is typically an exceedingly large quantity, in practice $P$ is usually truncated to its first $Q$ rows, with $Q$ being the maximum number of queries allowed; for our information-theoretic analysis in this paper, however, we do not consider the effect of truncation and hence all EPs of $P$ will be queried in the worst case.
As shown in [11, Sec. II], ORBGRAND and many other variants of GRAND can be represented by the following unified decoding rule, as long as no truncation of $P$ is applied, i.e., $Q=2^{N}$:
@@
The form of (2) facilitates information-theoretic analysis, as will be conducted in the following sections. A decoding error event occurs if $\hat{\mathsf{W}}\neq\mathsf{W}$, where $\hat{\mathsf{W}}$ is the decoder output, and we define the achievable rate of GRAND as follows.
@@
If we do not restrict the decoding rule to be (2) and apply ML decoding, then for our system model in Section II-A, according to Shannon’s channel coding theorem, the supremum of all achievable rates is the mutual information $I(\mathsf{X};\mathsf{Y})$ with $\mathsf{X}$ uniformly distributed over $\{+1,-1\}$. This mutual information thus serves as the ultimate performance limit for the considered system model, and is termed the symmetric capacity in the sequel.
@@ III Capacity-achieving Property of CDF-ORBGRAND
[…]
@@ III-A Main Result
Neither ORBGRAND nor CDF-ORBGRAND is ML decoding, and furthermore, their metrics, $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$, are $N$ correlated random variables due to ranking. Therefore, we cannot directly invoke existing information-theoretic formulas for evaluating their achievable rates. In this paper, we essentially establish the achievable rate of CDF-ORBGRAND as the generalized mutual information (GMI) of the considered decoding rule, following the general approach of mismatched decoding (see, e.g., [23, 24]). As revealed by the result below, although CDF-ORBGRAND is inherently mismatched, its achievable rate exactly coincides with the symmetric capacity. In other words, from an information-theoretic perspective, no rate loss is incurred when applying CDF-ORBGRAND.
@@
In our previous work[25], we have derived the GMI of ORBGRAND for the system model in Section II-A as
@@
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Psi\left(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\right)q^{+}(y)\mathrm{d}y$
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Psi\left(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\right)q^{-}(y)\mathrm{d}y\Bigg\}$ (4)
@@
We conduct some numerical experiments to compare (4) and (5), for antipodal signaling, i.e., BPSK modulation, over AWGN, additive white Laplacian noise (AWLN), and Rayleigh fading channels with receiver-side perfect channel state information (CSI). These channels can be written as
$\mathsf{Y}=\sqrt{P}\mathsf{H}\mathsf{X}+\mathsf{Z}:$
@@
  • Rayleigh fading channel with receiver-side perfect CSI, $\mathsf{H}\sim\mathcal{CN}(0,1)$ and $\mathsf{Z}\sim\mathcal{CN}(0,1)$: $\mathsf{T}=4\sqrt{P}\Re(\mathsf{H}^{*}\mathsf{Y})$. We point out that here $q^{\pm}(\cdot)$ should be evaluated for the augmented channel output $(y,h)$, i.e., including the receiver-side perfect CSI. [^4] We clarify that in the numerical experiments in [25], the GMI of ORBGRAND for this setup was incorrectly computed, due to an error in the program evaluating the integrals in (4). In the current paper, we rectify the error and provide updated numerical results. The program code can be found at: https://github.com/ustclizhuang/Simulation
@@ III-B Proof of Main Result
Since the terms in the summation in (6) involve ranks of channel reliabilities, they are correlated, and hence the standard GMI formula (see, e.g., [23, Eqn. (12)]) cannot be straightforwardly applied. Therefore we conduct the analysis directly from the first principle, akin to that in [11, 25] and dating back to [24, 26]. We calculate the ensemble average error probability of decoding, which, for the i.i.d. random coding ensemble in our system model, is equivalent to the error probability of decoding when message $w=1$ is transmitted.
Under the condition that the transmitted message is $w=1$, some asymptotic properties of the CDF-ORBGRAND decoding metric (6) are presented in the following lemmas, whose proofs are provided in the appendices.
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\right\}$
$=$ $\mathbb{E}\left[\ln\left(1+e^{\theta\cdot\bigl\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\bigr\rvert}\right)\right]-\ln 2.$ (9)
@@
$=\quad$ $\text{Pr}[\hat{\mathsf{W}}\neq 1|\mathcal{U}_{\epsilon}]\text{Pr}[\mathcal{U}_{\epsilon}]+\text{Pr}[\hat{\mathsf{W}}\neq 1,\mathcal{U}_{\epsilon}^{c}]$
@@
$\lim\limits_{N\to\infty}$ $\text{Pr}\Biggl[\mathsf{D}(1)\geq\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Biggr]=0.$ (11)
@@
On the other hand, from the decoding rule (2) together with the union bound, we obtain
@@
$\leq\quad$ $\text{Pr}\Biggl[\exists w^{\prime}\neq 1,\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Biggr]$
@@
Considering the probability in (12) conditioned upon $\underline{\mathsf{T}}$, and applying Chernoff bound, we obtain that for any $N$ and any $\theta<0$,
$-\frac{1}{N}\ln\text{Pr}\Biggl[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg|\underline{\mathsf{T}}\Biggr]$
$\geq\quad$ $\theta\cdot\Biggl[\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Biggr]$
$-\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\right\}.$ (13)
@@
$\text{Pr}\Biggl[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg|\underline{\mathsf{T}}\Biggr]$
@@
$\hskip 40.00006pt+\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Biggr]\Bigg\}.$ (14)